a ball is dropped onto a floor fom a height of 10m. if 20% of its initial energy is lost , then the height of bounce is (1)2m (2)4m (3)8m (4)6.4m

Grade:11

1 Answers

Vikas TU
7 years ago
mgh = 0.5mv^2
v = root(2gh) => root(2*g*10) = > 10root(2) m/s
K.E = 0.5m*200
80% of KE = >  0.5m*200*80/100 => 80m
0.5mv^2 = 80m
v^2 = 160
From third law of eqn.,
0^2 = 160 – 2gs
20s = 160
s = 8 m